# Decomposition of Borel graphs and cohomology

### v2

@@ 5 Proof of the theorems
First, we construct the standard Borel space $Y$ by adding new vertices on edges of $G$. Take a Borel subset $G^{+}\subset G$ so that for every $(x_{0},x_{1})\in G$, we have |$|\{(x_{0},x_{1}),(x_{1},x_{0})\}\cap G^{+}|=1$. By Lemma 3.11, we have
@@
$C^{\prime}\subset B_{G^{\prime}}\left(r;\lambda^{-1}(C)\right)\ \textup{ and }\ \overline{C^{\prime}}\subset B_{G^{\prime}}\left(r;\overline{\lambda^{-1}(C)}\right)$ (5.1)
@@
Now let $\xi\in Z_{H}$ be defined by $\xi(\cdot,y)=1_{\lambda^{-1}(C(y))\cap[y]_{H}}$ with $C(y)\in\mathcal{C}\cup\{\varnothing\}$. Define $\widehat{\xi}\in Z_{G^{\prime}}$ by $\widehat{\xi}(\cdot,y)=1_{\lambda^{-1}(C(y))}$. By Lemma 4.25 and claim (5.1), the function $\widehat{\xi}$ is represented by $\mathcal{C^{\prime}}$. This implies that $\xi=\widehat{\xi}|_{E_{H}}$ is represented by ${\mathcal{C}^{\prime}|_{H}}^{*}=\varnothing$. Hence $\xi$ is trivial in $\textup{H}^{1}(H,R_{H})$. ∎
The next lemma is also needed by a technical reason.
@@
for $10\leq i\leq n-$ 1. Suppose toward a contradiction that $y_{0}=y_{2n}$. Since $\gamma$ is injective, we can take $(x_{2i-1},x_{2i})\in E_{K}\setminus\Delta_{X}$ so that $(\gamma(x_{2i-1}),\gamma(x_{2i}))=(y_{2i-1},y_{2i})$ for $1\leq i\leq n$. We set $x_{0}=x_{2n}$. Then we have $(x_{2i},x_{2i+1})\in E_{H}\setminus\Delta_{X}$ for $10\leq i\leq n-$ 1 since $E_{H}=(\gamma\times\gamma)^{-1}(E_{H^{\prime}})$, which follows from that $\gamma$ is a quasi-isometry. This contradicts $G=H\ast K$. Hence $E_{H^{\prime}}$ and $E_{(\gamma\times\gamma)(K)}$ are freely intersecting.
@@
for every $11\leq i\leq n-$ 1. Set $G^{\prime}=H^{\prime}\ast(\gamma\times\gamma)(K)$. Then we have
@@
Now we write Theorem 1.3 again, and prove it.
@@
    If $\lambda_{k}:(X_{k},H_{k}^{\prime})\rightarrow(X_{k-1},H_{k-1})$ is a Borel quasi-isometric inverse of $\gamma_{k}$ with $\lambda_{k}\circ\gamma_{k}=\textup{id}_{X_{k-1}}$, then every element of $Z_{H_{k}}$ represented by $\lambda_{k}^{-1}{\left.\left(\mathcal{C}_{k}^{k-1}\right)\right|_{H_{k}}}^{*}$ is trivial in $\textup{H}^{1}(H_{k},R_{H_{k}})$.
By assumption (ix) on $(X_{k-1},G_{k-1})$ and condition (xii), the set $\textup{H}^{1}(H_{k},R_{H_{k}})$ is generated by the cutset $\bigcup_{i=k+1}^{n}\mathcal{C}_{i}^{k}$, where $\mathcal{C}_{i}^{k}=\lambda_{k}^{-1}{\left.\left(\mathcal{C}_{i}^{k-1}\right)\right|_{H_{k}}}^{*}$ is a Borel treeset of $H_{k}$. Now by Lemma 5.2 (applying to $(X,G)=(X_{k-1},G_{k-1})$, $H=H_{k-1}$, $K=T_{k-1}$ and $\gamma=\gamma_{k}:(X_{k-1},H_{k-1})\to(X_{k},H_{k}^{\prime})$), the equivalence relations $E_{H_{k}^{\prime}}$ and $E_{(\gamma_{k}\times\gamma_{k})(T_{k-1})}$ are freely intersecting, and the map
@@
In some sense, our proof above is rougher than Dunwoody’s proof of Theorem 1.2 (see also [DD, IV, Theorem 7.5]). One of the differences is that Dunwoody constructs a single treeset separating all ends of the group. By the argument in the proof of [Jar, Theorem 4.3], for any locally finite Borel graph $(X,G)$ and any integer $n\geq 1$, we can take a Borel treeset of $G$ with edge boundaries of size at most $n$ which generates the Boolean algebra on $X$ generated by all such cuts. However since we are concerned with the metric structure of Borel graphs, this machinery is not fit for our purpose. Therefore, instead of taking a single nice treeset, we divide the cutset into finitely many treesets and decompose the Borel graph several times. This might be impossible for general locally finite Borel graphs, and it is one of the reasons why we consider the case of uniformly bounded degrees.
5.2 Proof of Theorem 1.15
We write Theorem 1.15 for the reader’s convenience.
@@
Implication (i)$\Rightarrow$(ii) follows from Lemma 4.9. To show the converse, we assume $\textup{cd}_{R}(G)\leq 1$. Then $\textup{H}^{1}(G,R_{G})$ is finitely generated as a right $R_{G}$-module by Lemma 4.12. By Theorem 1.3, there exist a Borel graph $(Y,G^{\prime})$ with uniformly bounded degrees and an injective Borel quasi-isometry $\gamma:(X,G)\to(Y,G^{\prime})$ such that:
@@
Since a Borel quasi-isometry preserves the cohomological dimension by Lemma 4.14, we have $\textup{cd}_{R}(G^{\prime})\leq 1$, which implies $\textup{cd}_{R}(H)\leq 1$ by Lemma 4.13. Then the conditions $\textup{cd}_{R}(H)\leq 1$ and $\textup{H}^{1}(H,R_{H})=0$ imply that $\textup{cd}_{R}(H)=0$ by Lemma 4.10, and thus $E_{H}$ is uniformly finite by Lemma 4.11. Hence $H$ can by replaced by an Borel acyclic graph $T_{2}$ on $Y$ Lipschitz equivalently. Then the Borel graph $T^{\prime}=T_{1}\ast T_{2}$ is acycilc and Lipschitz equivalent to $G^{\prime}$. Note that $\gamma:(X,G)\to(Y,T^{\prime})$ is also a Borel quasi-isometry.
@@
Acknowledgments. The author would like to thank his advisor Yoshikata Kida for helpful comments and suggestions. He is also grateful to Anush Tserunyan for valuable conversations. This work was supported by JSPS KAKENHI Grant Number 23KJ0653, and the WINGS-FMSP program in the University of Tokyo.

### v3

@@ 5.1. Proof of Theorem A
First, we construct the standard Borel space $Y$ by adding new vertices on edges of $G$. Take a Borel subset $G^{+}\subset G$ so that for every $(x_{0},x_{1})\in G$, we have $|\{(x_{0},x_{1}),(x_{1},x_{0})\}\cap G^{+}|=1$. By Lemma 3.11, we have
@@
[…]
@@
Now let $\xi\in Z_{H}$ be defined by $\xi(\cdot,y)=1_{\lambda^{-1}(C(y))\cap[y]_{H}}$ with $C(y)\in\mathcal{C}\cup\{\varnothing\}$. Define $\widehat{\xi}\in Z_{G^{\prime}}$ by $\widehat{\xi}(\cdot,y)=1_{\lambda^{-1}(C(y))}$. By Lemma 4.25 and claim (5.1), the function $\widehat{\xi}$ is represented by $\mathcal{C^{\prime}}$. This implies that $\xi=\widehat{\xi}|_{E_{H}}$ is represented by $\mathcal{C}^{\prime}|_{H}$$^{*}$$\varnothing$. Hence $\xi$ is trivial in $\textup{H}^{1}(H,R_{H})$. ∎
We also need the following lemma.
@@
for $0\leq i\leq n-1$. Suppose toward a contradiction that $y_{0}=y_{2n}$. Since $\gamma$ is injective, we can take $(x_{2i-1},x_{2i})\in E_{K}\setminus\Delta_{X}$ so that $(\gamma(x_{2i-1}),\gamma(x_{2i}))=(y_{2i-1},y_{2i})$ for $1\leq i\leq n$. We set $x_{0}=x_{2n}$. Then we have $(x_{2i},x_{2i+1})\in E_{H}\setminus\Delta_{X}$ for $0\leq i\leq n-1$ since $E_{H}=(\gamma\times\gamma)^{-1}(E_{H^{\prime}})$, which follows from that $\gamma$ is a quasi-isometry. This contradicts $G=H\ast K$. Hence $E_{H^{\prime}}$ and $E_{(\gamma\times\gamma)(K)}$ are freely intersecting.
@@
for every $1\leq i\leq n-1$. Set $G^{\prime}=H^{\prime}\ast(\gamma\times\gamma)(K)$. Then we have
@@
Now we write Theorem A again, and prove it.
@@
    If $\lambda_{k}:(X_{k},H_{k}^{\prime})\rightarrow(X_{k-1},H_{k-1})$ is a Borel quasi-isometric inverse of $\gamma_{k}$ with $\lambda_{k}\circ\gamma_{k}=\textup{id}_{X_{k-1}}$, then every element of $Z_{H_{k}}$ represented by $\lambda_{k}^{-1}{{(\mathcal{C}_{k}^{k-1})}|_{H_{k}}}^{*}$ is trivial in $\textup{H}^{1}(H_{k},R_{H_{k}})$.
By assumption (ix) on $(X_{k-1},G_{k-1})$ and condition (xii), the set $\textup{H}^{1}(H_{k},R_{H_{k}})$ is generated by the cutset $\bigcup_{i=k+1}^{n}\mathcal{C}_{i}^{k}$, where $\mathcal{C}_{i}^{k}=\lambda_{k}^{-1}{(\mathcal{C}_{i}^{k-1})|_{H_{k}}}^{*}$ is a Borel treeset of $H_{k}$. Now by Lemma 5.2 (applying to $(X,G)=(X_{k-1},G_{k-1})$, $H=H_{k-1}$, $K=T_{k-1}$ and $\gamma=\gamma_{k}:(X_{k-1},H_{k-1})\to(X_{k},H_{k}^{\prime})$), the equivalence relations $E_{H_{k}^{\prime}}$ and $E_{(\gamma_{k}\times\gamma_{k})(T_{k-1})}$ are freely intersecting, and the map
@@
We remark about a feature of the above proof compared to Dunwoody’s proof of Theorem 1.2 (see also [DD89, IV, Theorem 7.5]). One of the differences is that they provide a concrete construction of a single treeset separating all ends of the group (which is invariant under the group action). A similar thing is possible in our setting. Indeed, by the argument in the proof of [Jar23, Theorem 4.3], for any locally finite Borel graph $(X,G)$ and any integer $n\geq 1$, there exists a Borel treeset of $G$ with edge boundaries of size at most $n$ which generates the Boolean algebra on $X$ generated by all such cuts. However since we are concerned with the metric (not the size) of the boundaries of cuts, this machinery is not fit for our purpose. Therefore, instead of taking a single nice treeset, we divide the cutset into finitely many treesets and decompose the Borel graph several times. This might be impossible for general locally finite Borel graphs, and it is one of the reasons why we restrict our attention to the case of uniformly bounded degrees.
5.2. Proof of Theorem B
We recall the statement of Theorem B:
@@ 5.2. Proof of Theorem B
[…]
@@
Since a Borel quasi-isometry preserves the cohomological dimension by Lemma 4.14, we have $\textup{cd}_{R}(G^{\prime})\leq 1$, which implies $\textup{cd}_{R}(H)\leq 1$ by Lemma 4.13. Then the conditions $\textup{cd}_{R}(H)\leq 1$ and $\textup{H}^{1}(H,R_{H})=0$ imply that $\textup{cd}_{R}(H)=0$ by Lemma 4.10, and thus $E_{H}$ is uniformly finite by Lemma 4.11. Hence $H$ can be replaced by an Borel acyclic graph $T_{2}$ on $Y$ Lipschitz equivalently. Then the Borel graph $T^{\prime}=T_{1}\ast T_{2}$ is acycilc and Lipschitz equivalent to $G^{\prime}$. Note that $\gamma:(X,G)\to(Y,T^{\prime})$ is also a Borel quasi-isometry.
@@
Acknowledgments
I would like to thank my advisor, Professor Yoshikata Kida, for his continuous mentorship throughout my master’s and Ph.D. studies. This thesis was completed thanks to his helpful comments and suggestions. I am also grateful to Professor Anush Tserunyan for her warm hospitality during my visit to McGill University and for many valuable discussions. This work was supported by JSPS KAKENHI Grant Number 23KJ0653, and the WINGS-FMSP program in the University of Tokyo.